### v1

# Secret Breach Detection in Source Code with Large Language Models

Abstract
Background: Leaking sensitive information—such as API keys, tokens, and credentials—in source code remains a persistent security threat. Traditional regex and entropy-based tools often generate high false positives due to limited contextual understanding. Aims: This work aims to enhance secret detection in source code using large language models (LLMs), reducing false positives while maintaining high recall. We also evaluate the feasibility of using fine-tuned, smaller models for local deployment. Method: We propose a hybrid approach combining regex-based candidate extraction with LLM-based classification. We evaluate pre-trained and fine-tuned variants of various Large Language Models on a benchmark dataset from 818 GitHub repositories. Various prompting strategies and efficient fine-tuning methods are employed for both binary and multiclass classification. Results: The fine-tuned LLaMA-3.1 8B model achieved an F1-score of 0.9852 in binary classification, outperforming regex-only baselines. For multiclass classification, Mistral-7B reached 0.982 accuracy. Fine-tuning significantly improved performance across all models. Conclusions: Fine-tuned LLMs offer an effective and scalable solution for secret detection, greatly reducing false positives. Open-source models provide a practical alternative to commercial APIs, enabling secure and cost-efficient deployment in development workflows.
Index Terms:
Source Code, Secret. Large Language Model, Regular Expression
I Introduction
Secrets play a crucial role in software development by serving as a digital authentication mechanism that protects access to sensitive systems and services. These secrets take many forms, such as API tokens, OAuth credentials, encryption keys (notably RSA), TLS/SSL certificates, and authentication details like usernames and passwords [1]. However, as modern applications rely on interconnected services, secrets can unintentionally spread across different components, amplifying the risk of exposure [2]. Figure 1 shows an example of secret breach in source code. The JavaScript code above illustrates the use of the fetch API to perform an authenticated HTTP GET request. The authorization is handled via a Bearer token included in the request header. The response is parsed as JSON and logged to the console. Any errors encountered during the request are captured and reported.
Figure 1: Example of a secret breach in source code: An API key is hardcoded and exposed, posing a security risk if the code is shared or pushed to a public repository.
Platforms like GitHub and Bitbucket have become essential for developers to collaborate and share code. Unfortunately, this also means that secrets are often inadvertently exposed within repositories. GitGuardian’s annual report on secret leaks highlights a sharp increase in such incidents since 2020. In 2022 alone, 10 million hardcoded secrets were detected—marking a 67% rise compared to 2021—while scanned commits surged from 860 million in 2021 to over 1 billion in 2022 [3]. Numerous reports have documented cases where attackers exploited exposed API keys from public repositories to access free services [4]. A study by Sophos revealed that compromised credentials were responsible for half of all recorded cyberattacks in the first half of 2023 [5]. Despite efforts by cloud providers like Amazon and Google to detect and revoke leaked keys, version control systems (VCS) offer minimal protections. Beyond source code, secrets may also be unintentionally leaked through configuration files, test scripts, documentation, dependency files, and logs [2].
The 2024 report [6] also identified a significant rise in GenAI-related secret leaks, most notably a 1212 times increase in leaked OpenAI API keys, making them the highest growth detector class. These secrets were leaked at an alarming rate of 46,000 per month on average. Equally concerning is the rise of “zombie leaks”—secrets that remain valid even after repositories are deleted or made private. Such secrets are especially dangerous as they linger without visibility, creating long-term vulnerabilities.
A variety of tools [7] exist to identify vulnerabilities during code commits, including TruffleHog [8], Gitrob [9], ggshield [10], and Gitleaks [11]. These tools typically rely on regex-based [12] pattern matching and entropy analysis. For instance, Gitleaks and ggshield use predefined regex patterns to detect hardcoded secrets, whereas TruffleHog identifies high-entropy strings [13]. However, such techniques often yield high false positive rates [14], making them less effective in noisy environments.
To address these limitations, Meli et al. [15] conducted a comprehensive study on GitHub repositories, combining pattern matching with entropy-based filtering via regex. Their approach successfully detected secrets in over 100,000 repositories with an accuracy of 99.29%. Further advancements were made by Sinha et al. [16], who incorporated keyword searches and program slicing techniques to improve both recall and precision in identifying API keys and similar credentials. Meanwhile, Saha et al. [17] tackled false positives by training machine learning models on datasets retrieved through the Git REST API [18], achieving an F1-score of 86.7% using a Soft-Voting Classifier. Feng et al. [19] further refined secret detection by applying deep neural networks to reduce false positives in textual password detection. To mitigate the risks associated with secret leaks, Sinha et al. [16] suggested various preventive measures, including the use of tools like git-crypt [20], which encrypts sensitive data in Git repositories. Wahab et al. [21] used transformer models like BERT [22] to detect secrets in Github issue reports. Additionally, Basak et al. [23] compiled 24 best practices for effective secret management, offering structured guidelines to minimize exposure risks
In this work, we propose a novel approach to automate secret detection in source code by leveraging large language models (LLMs). Our methodology begins with regex-based candidate string extraction, after which an LLM classifies true secrets, significantly reducing false positives. We experiment with various strategies, including zero-shot, one-shot, few-shot learning, and zero-shot inference after fine-tuning. Our fine-tuned model achieves an F1-score of 0.9852 on the test dataset, outperforming regex-based and conventional machine learning approaches.
Our key contributions are as follows:
  • We develop a hybrid methodology that first uses regex-based candidate extraction and then refines classification using LLMs, reducing false positives while maintaining high recall.
  • We test multiple LLM-based detection strategies, including zero-shot, one-shot, few-shot learning, and fine-tuned inference, to assess their impact on precision and recall.
  • We extend the detection approach to multiclass classification, categorizing secrets into types such as Private Keys, API Keys, Authentication Tokens, and more, significantly enhancing the granularity of detection.
To this end, we aim to address the following research questions:
  • RQ1: How effective are large language models in detecting secrets?
  • RQ2: How effective are smaller models for local server deployment?
The rest of this paper is organized as follows. Section II-A describes the methodology, including data collection, preprocessing, and secret detection using large language models. Section II-B presents our LLM-based detection framework, covering prompting strategies and performance across different models. Section II-C discusses the use of fine-tuned smaller LLMs for local deployment. Section III provides a detailed discussion on evaluation insights, limitations, and impact factors. Section IV reviews related work. Section V outlines potential threats to validity, and finally, Section VI concludes the paper and discusses future directions.
Replication Package. Our code and data are shared at https://doi.org/10.5281/zenodo.15273217
II Methodology
II-A Data Collection and Preprocesssing
II-A1 Data Collection
Our dataset is derived from SecretBench [24], a publicly available benchmark dataset containing 97,479 candidate secrets extracted from 818 public GitHub repositories. Among these, 15,084 are labeled as true secrets, manually verified by analyzing both the secret strings and their surrounding code context. The dataset spans 49 programming languages and 311 different file types, covering various forms of secrets such as API keys, OAuth tokens, cryptographic private keys, and database credentials.
The original SecretBench dataset was constructed using 761 predefined regular expressions (regex)[25] to identify potential secrets from GitHub repositories. These regex patterns targeted a diverse range of secret types, including high-entropy strings, access tokens, and password patterns. Each detected candidate was manually labeled to distinguish between actual secrets and false positives, such as dummy credentials or placeholders. The dataset also includes metadata such as file paths, commit IDs, entropy scores, and programming language information.
II-A2 Preprocessing
To prepare the dataset for training and evaluation, we performed multiple preprocessing steps:
  1. 1.
    File Content Extraction – Each candidate secret’s corresponding source file was retrieved from the Files.zip dataset. File contents were read while handling encoding variations, with unreadable characters removed to ensure consistency.
  2. 2.
    Dataset Balancing – Since the dataset is highly imbalanced, we curated a subset for training and evaluation:
    • 30,000 non-secret samples were randomly selected.
    • 15,000 true secret samples were retained.
For multiclass classification, we only took the positive samples from the dataset. The dataset distribution is mentioned in table I.
TABLE I: Multiclass Secret Distribution (15,084 samples)
Secret Type Count
Private Key 5,791
API Key and Secret 4,521
Authentication Key and Token 3,567
Other 524
Generic Secret 333
Database and Server URL 162
Password 150
Username 27
To train our LLMs on multiclass secret detection, we split the labeled dataset as follows:
  • Training Set: 9,000 samples
  • Evaluation Set: 3,000 samples
  • Test Set: 3,000 samples
We followed the same preprocessing steps as in binary classification.
Figure 2 illustrates the overall workflow of our proposed secret detection system. The process begins with the extraction of candidate strings using a set of predefined regular expressions. For each candidate, we construct a contextual window from the surrounding source code to provide the necessary semantic cues. These context-enriched inputs are then classified using either pre-trained or fine-tuned large language models (LLMs). The figure serves as a high-level guide to the methodology, with each component discussed in detail in the following sub-sections.
Figure 2: Workflow for secret detection in source code. Potential candidate strings are extracted using 761 regular expressions from a .csv file and located within the source code. A 200-character context window around each candidate is then created. This context is either fed directly into a pretrained LLM for zero-shot, one-shot, or few-shot classification, or passed through a fine-tuned LLM (trained on the SecretBench dataset) for downstream zero-shot classification.
II-B Secret Detection Using Pre-trained Large Language Models (RQ1)
II-B1 Motivation
Pre-trained large language models (LLMs) have shown remarkable performance have demonstrated exceptional capabilities in text classification [26] and contextual understanding [27], making them well-suited for secret detection. Their ability to understand code semantics [28] without extensive task-specific fine-tuning makes them attractive for security-related applications as well. Motivated by this, we investigate the inherent capability of LLMs to identify secrets embedded in source code. Specifically, we explore how well these models can distinguish true secrets from non-sensitive strings using contextual understanding, even without extensive retraining.
II-B2 Approach
In this study, we evaluated GPT-4o [29] and Deepseek-V3 [30] on detecting secrets as and experimented with different prompting strategies such as zero-shot, one-shot, and few-shot learning.
Context Extraction and Data Preprocessing: Secret detection often requires an understanding of the surrounding context [31] [32] in which a candidate string appears. To provide sufficient contextual information, we extract a 200-character window around each candidate secret from its source code. This allows the model to better differentiate between sensitive and non-sensitive patterns. Moreover, since source code files can span thousands of lines, feeding the entire file into a model is computationally expensive and often impractical.
Each data point consists of:
  • Candidate String: The potential secret identified from the dataset.
  • Code Snippet: The extracted surrounding code providing contextual information.
  • Label: A binary classification label indicating whether the candidate is a Secret or Non-sensitive.
The extracted data is then formatted into structured natural language prompts to facilitate LLM-based classification.
Prompt Engineering: The interaction with large language models (LLMs) is primarily conducted through prompt engineering, where a carefully crafted textual input is provided to the model to perform a specific task. In our case, we prompt the model to classify whether a given code snippet contains a secret or not. This is achieved by presenting the model with candidate strings and their surrounding context in source code. We explore several prompting strategies, including zero-shot, one-shot, and few-shot prompting, inspired by prior work on LLM-driven generation tasks such as test case creation and code understanding.
Zero-Shot Prompting
In the zero-shot setting, the model receives only the candidate string and its surrounding code snippet, without any labeled examples or instructions. It relies solely on its pre-trained knowledge to determine whether the string is a secret. This setting serves as a baseline to assess the model’s inherent reasoning ability. The prompt format used is shown in Figure 3.
Figure 3: Prompt for classifying a candidate string as ”Secret” or ”Non-sensitive” from a code snippet using pre-trained LLMs.
One-Shot Prompting
In one-shot prompting, a single labeled example with its code context is provided before the prediction task. This example demonstrates the classification process and guides the model’s decision-making. We select a general yet representative sample capturing common secret patterns.
Few-Shot Prompting
Few-shot prompting [33] extends the one-shot setting by including multiple labeled examples. These cover varied secret and non-secret cases, offering richer context. This setup improves model performance, especially on ambiguous inputs, by allowing it to learn from analogical reasoning.
Parameter Settings: To control the behavior of LLM outputs during inference, we adjust key generation parameters. The temperature parameter controls the randomness of the output distribution. Lower values make the model more deterministic, while higher values introduce more variability. We use a low temperature value (set to 0.0) to reduce randomness and ensure deterministic outputs.
Evaluation Data: For evaluation, we randomly selected a subset of 3000 (1500 secret + 1500 non-secret) data points from our dataset.
Evaluation Metrics: To assess the performance of our LLM-based secret detection models, we employ standard classification evaluation metrics: precision, recall, and F1-score. These metrics provide insights into how well the model distinguishes between secrets and non-sensitive strings.
  • Precision: Precision measures how many of the instances predicted as secrets are actually true secrets. It reflects the proportion of true positives among all predicted positives. A high precision score indicates fewer false alarms, reducing unnecessary security alerts.
  • Recall: Recall measures how many actual secrets were correctly identified by the model. It captures the proportion of true positives among all actual positives. A high recall score ensures that most secrets are identified, reducing the risk of sensitive data exposure.
  • F1-Score: The F1-score is the harmonic mean of precision and recall. It provides a balanced performance measure, especially useful when both false positives and false negatives are important to minimize.
Justification for Using F1-Score Since our dataset is balanced (12,000 secrets and 12,000 non-secrets), accuracy serves as a reasonable metric during training. However, in real-world scenarios, secret detection often requires balancing precision and recall. A model focused too heavily on precision may miss actual secrets, while one optimized for recall may trigger too many false alarms. To address this, we use the F1-score as our primary evaluation metric. It provides a balanced measure that minimizes both false positives and false negatives, which is crucial in security workflows where over-detection causes alert fatigue and under-detection can lead to serious breaches.
TABLE II: Evaluation Metrics for zero-shot learning on pretrained llms
Model Precision Recall F1-Score
DeepSeek-V3 0.7810 0.9033 0.8377
Gpt-4o 0.8233 0.8633 0.8428
TABLE III: Evaluation Metrics for one-shot learning on pretrained llms
Model Precision Recall F1-Score
DeepSeek-V3 0.7761 0.9913 0.8716
Gpt-4o 0.8623 0.9813 0.9180
TABLE IV: Evaluation Metrics for few-shot learning on pretrained llms
Model Precision Recall F1-Score
DeepSeek-V3 0.7926 0.9887 0.8799
Gpt-4o 0.9329 0.9465 0.9392
II-B3 Result
The results of DeepSeek and Gpt-4o, as presented in Table II, Table III, and Table IV, demonstrate that Gpt-4o consistently outperforms DeepSeek across all evaluation metrics—Precision, Recall, and F1-Score—in zero-shot, one-shot, and few-shot settings.
  • Zero-Shot: In the zero-shot setting, Gpt-4o achieves a Precision of 82.33%, Recall of 86.33%, and an F1-Score of 84.28%, outperforming DeepSeek across all metrics. This suggests that Gpt-4o is better equipped to utilize its pre-trained knowledge to identify sensitive strings from code snippets even without any labeled examples.
  • One-Shot: With the addition of a single labeled example, Gpt-4o shows substantial improvement—particularly in Precision, which rises to 86.23%, and F1-Score, which reaches 91.80%. Compared to DeepSeek, Gpt-4o demonstrates stronger generalization from minimal guidance.
  • Few-Shot: In the few-shot setting, Gpt-4o achieves its highest scores: 93.29% Precision, 94.65% Recall, and 93.92% F1-Score. These results surpass those of DeepSeek and indicate that Gpt-4o can effectively leverage multiple examples to distinguish secrets from non-sensitive content with exceptional accuracy and coverage.
Overall, Gpt-4o demonstrates stronger performance than DeepSeek across all prompting configurations, making it a more effective and reliable choice for secret detection tasks in source code.
II-B4 Multiclass Classification
As mentioned in Section II-A we extended our approach to multiclass classification. We evaluated using 3000 test samples from the partition mentioned in the dataset. Despite their general language understanding capabilities, these models exhibit significantly lower performance in secret type classification compared to our finetuned counterparts.
Table V presents the overall classification performance, with GPT-4o achieving a weighted F1-score of 0.8108. The weighted F1-score was chosen as the evaluation metric to account for class imbalance in the dataset.
TABLE V: Overall Multiclass Classification Accuracy for Pretrained Models
Model F1-score(weighted)
DeepSeek-V3 0.7338
GPT-4o 0.8108
Table VI presents the detailed classification report. While classes like Private Key and API Key and Secret achieve relatively high precision and recall, other categories such as Generic Secret, Other, and Username show poor performance, suggesting that zero-shot or few-shot capabilities are insufficient for nuanced secret-type recognition.
TABLE VI: Multiclass Classification Report for GPT-4o
Category Precision Recall F1-Score
Private Key 0.9897 0.9411 0.9648
API Key and Secret 0.7491 0.6667 0.7055
Authentication Key and Token 0.9009 0.8276 0.8627
Other 0.2373 0.1429 0.1783
Generic Secret 0.0844 0.4407 0.1417
Database and Server URL 0.5581 0.6857 0.6154
Password 0.8485 0.7179 0.7778
Username 0.0000 0.0000 0.0000
Weighted F1-score 0.8108
These findings highlight the limitations of relying solely on pretrained models for domain-specific classification tasks like secret type detection, where fine-tuning significantly improves both accuracy and category-level performance.
RQ1 Summary: This study investigates the ability of pre-trained large language models (LLMs), specifically GPT-4o and DeepSeek-V3, to detect secrets in source code without extensive fine-tuning. By using context windows of 200 characters and structured prompts, the models were evaluated in zero-shot, one-shot, and few-shot learning settings. GPT-4o consistently outperformed DeepSeek across all prompting strategies, achieving the highest F1-score of 93.92% in the few-shot setting. Although effective at binary classification, the models did not perform remarkably well in multiclass secret-type classification, highlighting the need for fine-tuning to achieve reliable performance in nuanced tasks.
II-C Secret Detection with Smaller Fine-tuned Models (RQ2)
II-C1 Motivation
Many organizations are reluctant to adopt third-party large language models (LLMs) for software development tasks due to internal compliance policies, data privacy concerns, and the high costs associated with commercial APIs. To address these issues, we investigate the feasibility of using smaller, open-source models that can be fine-tuned and deployed on local servers. Models such as LLaMA [34], Gemma [35], CodeLLaMA [36], and DeepSeek offer cost-effective alternatives with promising capabilities. When fine-tuned with high-quality data, these models often achieve performance levels comparable to larger general-purpose LLMs, while consuming significantly less memory and compute resources. This makes them ideal for deployment in constrained environments. Moreover, locally hosted models eliminate usage-based costs and ensure complete control over sensitive and proprietary data, enhancing both security and compliance.
II-C2 Approach
We fine-tuned several compact models— LLaMA-3.1-8B, Mistral-7B, CodeLLaMA-7B, DeepSeek-7B, and Gemma-7B to detect secrets. These models were selected for their strong performance on complex language tasks, efficient inference capabilities, and suitability for long-sequence processing, all while maintaining a low computational footprint. In addition to evaluating the fine-tuned models by zero-shot prompting on them [37], we also assessed the raw, pre-trained versions to understand the benefits of fine-tuning. For the baseline comparison, we conducted zero-shot, one-shot, and few-shot prompting experiments on the pre-trained models without any task-specific training.
Context Extraction and Data Preprocessing: We used the same context window approach like section III-B.
Prompt Engineering: We used the same prompt format outlined in Section III-B for the fine-tuned models. We applied zero shot, one-shot and few-shot learning in raw pre-trained models without any task-specific training and applied zero-shot learning in fine-tuned models.
Parameter Settings: For fine-tuning our models, we employed QLoRA [38] with 4-bit precision to minimize memory consumption while preserving model effectiveness. Specifically, we used NF4 (Normal Float 4-bit) quantization with FP16 computation, and disabled double quantization to streamline training. Fine-tuning was carried out on transformer attention layers, enabling the models to learn domain-specific patterns crucial for accurate secret classification, even on resource-constrained hardware.
We adopted parameter-efficient fine-tuning (PEFT) [39] using the LoRA framework, which allows selective updating of model parameters. We set the LoRA rank to 64 and scaling factor $\alpha$ to 16. No dropout or bias modifications were used, ensuring faster training and reduced overhead. This setup facilitates robust performance while requiring only a fraction of the resources compared to full fine-tuning.
The models were fine-tuned using the Paged AdamW (32-bit) optimizer with a learning rate of 2e-4. This optimizer was chosen for its memory-efficient paging mechanism, which enables training large models on limited GPU resources without compromising performance. Training was performed for 7 epochs with a batch size of 1 per device and 8 gradient accumulation steps. A cosine decay learning rate scheduler was used, with a warm-up ratio of 3%.
II-C3 Evaluation Data
Due to the significant class imbalance in the original dataset (15,084 true secrets out of 97,479 candidates), we curated a balanced and representative subset for model training and evaluation. Specifically, we randomly selected 30,000 non-secret samples and retained all 15,000 true secrets, forming a pool of 45,000 labeled examples.
From this curated pool, we created four distinct subsets to support both balanced and imbalanced training scenarios, as well as model evaluation:
  • Balanced Train Set: 12,000 true secrets and 12,000 non-secrets (24,000 total)
  • Imbalanced Train Set: 3,750 true secrets and 20,250 non-secrets (24,000 total)
  • Validation Set: 1,500 true secrets and 1,500 non-secrets (3,000 total)
  • Test Set: 1,500 true secrets and 1,500 non-secrets (3,000 total, same set as Section II-B)
Evaluation Metrics: Similar to Section II-B, we used those evaluation metrics to evaluate the detected secrets.
TABLE VII: Evaluation Metrics for Zero-Shot Prompting (Raw Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.4105 0.4283 0.4054
Gemma-7B 0.5501 0.5003 0.3353
LLaMA-3.1 8B 0.5412 0.5130 0.4125
Mistral-7B 0.5033 0.5033 0.5029
CodeLLaMA-7B 0.5453 0.5383 0.5207
TABLE VIII: Evaluation Metrics for One-Shot Prompting (Raw Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.6864 0.5743 0.4990
Gemma-7B 0.6741 0.6610 0.6545
LLaMA-3.1-8B 0.7622 0.6920 0.6699
Mistral-7B 0.6848 0.5720 0.4949
CodeLLaMA-7B 0.6741 0.6610 0.6545
TABLE IX: Evaluation Metrics for Few-Shot Prompting (Raw Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.5216 0.5127 0.4563
Gemma-7B 0.7286 0.7187 0.7156
LLaMA-3.1-8B 0.7913 0.7913 0.7913
Mistral-7B 0.7678 0.7630 0.7619
CodeLLaMA-7B 0.7286 0.7187 0.7156
TABLE X: Evaluation Metrics for Zero-Shot Prompting (Fine-Tuned Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.9503 0.9473 0.9472
Gemma-7B 0.9756 0.9753 0.9753
LLaMA-3.1-8B 0.9861 0.9853 0.9852
Mistral-7B 0.9815 0.9825 0.9819
CodeLLaMA-7B 0.9716 0.9709 0.9707
II-C4 Result
The performance of all five LLMs—DeepSeek, Gemma, LLaMA-3.1-8B, Mistral, and CodeLLaMA—under different prompting settings without fine-tuning is summarized in Table VII, Table VIII, and Table IX. Table X compares their performance in the zero-shot setting after fine-tuning. Across all scenarios, the results clearly show that fine-tuning significantly boosts performance. Among the pre-trained models, LLaMA-3.1 8B consistently achieves the best F1-scores in all prompting conditions.
  • Zero-Shot (Raw): As seen in Table VII, model performance is relatively low, with F1-scores ranging from 0.33 to 0.52. CodeLLaMA-7B performs the best in this setting with an F1-score of 0.5207.
  • One-Shot (Raw): Table VIII shows a noticeable improvement in performance. LLaMA-3.1-8B achieves the highest F1-score of 0.6699, followed by Gemma and CodeLLaMA.
  • Few-Shot (Raw): With more contextual examples, performance improves further (Table IX). LLaMA-3.1 8B again leads with an F1-score of 0.7913, followed by Mistral-7B and Gemma-7B.
  • Zero-Shot (Fine-Tuned): As shown in Table X, fine-tuning significantly enhances model effectiveness. LLaMA-3.1-8B attains the top F1-score of 0.9852, followed by Gemma and CodeLLaMA. Even DeepSeek-7B, which had low performance in the raw setting, achieves an impressive 0.9472 F1-score after fine-tuning.
The confusion matrix in Figure 4 illustrates the performance of the fine-tuned LLaMA-3.1-8B model under the zero-shot prompting in fine-tuned model. The model correctly identified 1,496 out of 1,500 negative (non-secret) instances and 1,463 out of 1,500 positive (secret) instances. Only 4 false positives and 37 false negatives were observed, indicating high precision and recall. These results are consistent with the high F1-score (0.9852) observed in Table X, affirming the model’s strong discriminative ability after fine-tuning. The near-diagonal dominance of the matrix further emphasizes the model’s robustness in accurately separating secret and non-secret examples.
Figure 4: Confusion Matrix for Zero-Shot prompting on fine-tuned LLaMA-3.1-8B model
Overall, these results demonstrate that fine-tuning and prompt-based learning significantly enhance the secret classification capabilities of LLMs, with LLaMA-3.1-8B consistently delivering top-tier performance.
II-C5 Multi class classification
To overcome the limitations of pretrained LLMs, we fine-tuned smaller language models on 9000 labeled positive samples, specifically curated for the secret classification task. These models were then evaluated on the same 3000-sample test set used for the pretrained models to ensure a fair comparison.
TABLE XI: Overall Multiclass Classification Accuracy
Model F1-Score(Weighted)
Deepseek-7B 0.9584
Gemma-7B 0.9494
Codellama-7B 0.9625
LLaMA-3.1-8B 0.980
Mistral-7B 0.982
TABLE XII: Multiclass Classification Report for Mistral-7B
Category Precision Recall F1-Score
Private Key 0.9875 0.9893 0.9884
API Key and Secret 0.9879 0.9760 0.9819
Authentication Key and Token 0.9972 0.9945 0.9959
Other 0.8788 0.8878 0.8832
Generic Secret 0.8060 0.9153 0.8571
Database and Server URL 1.0000 1.0000 1.0000
Password 0.9512 1.0000 0.9750
Username 1.0000 1.0000 1.0000
Overall Accuracy 0.9820
Table XII presents the classification reports for Mistral model. Figure 5 presents the confusion matrix for the Mistral-7B model, providing insights into its classification performance across different secret types. The model performs well on high-frequency categories, such as Private Key (1109 correct), API Key and Secret (896 correct), and Authentication Key and Token (721 correct), indicating a strong ability to distinguish these secret types.
Figure 5: Confusion Matrix for Mistral-7B fine tuned for multi class classification
While overall accuracy and weighted f1-scores are high, some misclassifications are observed. API Key and Secret secrets are occasionally predicted as Other or Authentication Key and Token, suggesting feature overlap. Similarly, the Other category exhibits greater misclassifications, reflecting its diverse nature. Lower-frequency categories, such as Generic Secret, Database and Server URL, and Password, are correctly classified in most cases, but minor errors remain.
RQ2 Summary: We evaluated the effectiveness of fine-tuned open-source LLMs (e.g., LLaMA-3.1 8B, Mistral, CodeLLaMA, DeepSeek, and Gemma) for secret detection using both binary and multiclass classification. Fine-tuning significantly improved performance, with LLaMA-3.1 8B achieving the highest F1 score (0.9852) in zero-shot settings. In multiclass classification, both LLaMA-3.1 8B and Mistral-7B achieved high accuracy (98%+), with Mistral showing strong category-wise precision and recall. These results affirm that compact, fine-tuned models are highly effective and resource-efficient alternatives to commercial LLMs for sensitive software engineering tasks like secret detection.
III Discussion
III-A Analysis of Detected Secrets
To evaluate the effectiveness of the fine-tuned LLaMA model in binary classification, we analyze its ability to distinguish between true secrets and non-secret content.
True Positives.
The model successfully identified sensitive information, such as an API key embedded in an HTTP request and a cryptographic private key used for decryption. These examples demonstrate the model’s robustness in recognizing structured secrets that adhere to common patterns. For instance, the API key (sk_test_4eC39HqLyjWDarjtT1zdp7dc) within a request to Stripe’s API was correctly classified as a secret, as was the cryptographic private key in a decryption function. This suggests that the model has effectively learned to detect well-structured secrets present in programming environments.
1import requests
2
3API_KEY = "sk_test_4eC39HqLyjWDarjtT1zdp7dc"
4
5response = requests.get(f"https://api.stripe.com/v1/charges", headers={{
6 "Authorization": f"Bearer API_KEY"
7}})
8print(response.json())
False Positives.
The model correctly classified non-secret content as non-sensitive, avoiding false positives in cases where misleading patterns could exist. One such example is a login function containing hardcoded credentials ("user123" and "password"). While these may appear to be sensitive at first glance, they lack complexity and randomness typical of actual secrets. Similarly, an obfuscated API key ("xxxxxxxxxx") was correctly ignored, likely due to the model recognizing it as a placeholder rather than an actual sensitive credential.
1def login_user(username, password):
2 if username == "user123" and password == "password":
3 print("Login successful")
4 else:
5 print("Login failed")
1import requests
2
3API_KEY = "xxxxxxxxxx"
4
5response = requests.get(f"https://api.stripe.com/v1/charges", headers={{
6 "Authorization": f"Bearer API_KEY"
7}})
8print(response.json())
These results highlight the model’s strong performance in distinguishing true secrets from false alarms. However, future improvements could focus on refining detection criteria to minimize edge cases where synthetic or low-entropy secrets might be misclassified.
III-B Comparison with Existing Tools
Regex-based tools such as TruffleHog are effective for identifying candidate secret strings but suffer from high false positive rates due to a lack of contextual understanding. Since our dataset is constructed from regex-extracted candidates, many of which are false positives, it inherently reflects this limitation. While they exhibit high recall, they suffer from low precision due to an inability to disambiguate true secrets from benign strings.
The best-performing fine-tuned LLaMA-3.1 8B model achieved an F1-score of 0.9852 on our subset of 3,000 samples (1,500 secrets and 1,500 non-secrets). In contrast, a regex-only tool would flag all candidates as secrets, yielding a theoretical precision of just 50% in this balanced setting.
Our LLM-based classifier thus significantly improves precision while maintaining high recall. By leveraging the contextual understanding of LLMs, we reduce noise introduced by regex and filter out false positives more effectively. This hybrid approach combines the broad detection capability of regex with the discriminative power of LLMs, making it more practical for integration into modern development workflows.
III-C False Negatives
Despite the strong performance of our fine-tuned models, a limited number of false negatives were observed during evaluation. Specifically, 37 instances were incorrectly classified as non-sensitive. A detailed manual analysis revealed several contributing factors behind these misclassifications.
One source of error was inconsistencies between our annotation framework and the labeling conventions of the SecretBench dataset [24]. For example, the dataset labels certain public keys—such as consensus public key—as sensitive, while our model and prompt design treated such keys as non-sensitive based on commonly accepted definitions of sensitive information. As a result, some predictions were considered false negatives relative to the dataset, although they were consistent with our defined classification criteria. A similar discrepancy was found in cases such as pk_test_..., which our model interpreted as a benign public key based on its naming pattern and prior examples. However, the dataset labeled it as sensitive, potentially reflecting real-world risks associated with leaked test tokens, even when such keys are publicly documented to be accessible from the client side (e.g., Stripe documentation). These examples highlight the ambiguity that can exist in the classification of security-relevant identifiers.
Another contributing factor was the model’s limited capacity to accurately detect secrets when the surrounding context was insufficient. This limitation was particularly evident in cases involving long-format secrets, such as RSA private keys, which often exceed the default input length. To mitigate this, we increased the model’s maximum input sequence length from 512 to 1024 tokens. This modification resulted in a measurable improvement, reducing the number of false negatives by enabling the model to process a more complete context for each candidate.
Additionally, we observed that very small number of false negatives arose due to subtle variations in context. Identical candidate strings appeared in different code environments and received inconsistent classifications. This behavior suggests that the model may be sensitive to minor contextual cues, leading to inconsistent decision-making. This represents a potential area for future improvement, particularly in enhancing the model’s robustness to contextual variability.
Overall, we acknowledge that the presence of a small number of false negatives reflects a trade-off aimed at significantly reducing false positives. By aligning the model’s predictions with a clearly defined set of criteria, we prioritized precision in sensitive data detection. In applications where minimizing false alarms is essential to maintaining developer trust and usability, we believe this trade-off is both justified and desirable.
III-D Impact of Imbalanced Training Distribution
In addition to balanced training, we also fine-tuned our models using an imbalanced training set that better reflects the real-world distribution of secrets in source code. Specifically, we used a training dataset comprising 3,750 true secrets and 20,250 non-secret samples, maintaining the total size at 24,000 instances. This distribution mirrors the actual ratio observed in the SecretBench dataset, where only 15,084 out of 97,479 samples are true secrets.
Despite the skewed class proportions, the performance of the fine-tuned models remained robust. As shown in Table XIII, our best-performing model, LLaMA-3.1-8B, achieved a precision of 0.9742, recall of 0.9730, and an F1-score of 0.9730 on a balanced evaluation set. When trained on the imbalanced dataset, these scores experienced only a slight drop (1–3% across most metrics), indicating that LLMs retain their contextual reasoning capabilities and strong classification performance even under class imbalance.
This resilience can be attributed to the expressive power of LLMs in understanding code semantics and distinguishing true secrets from non-sensitive tokens based on context. Unlike traditional classifiers that may suffer significantly in imbalanced scenarios, LLMs leverage large-scale pretraining and fine-grained token-level understanding to preserve high recall without compromising much on precision.
Our findings confirm that even when trained on datasets resembling real-world secret distributions, fine-tuned LLMs can maintain excellent detection capabilities. This makes them especially suitable for deployment in practical settings, where secrets are rare but their detection is critical.
TABLE XIII: Evaluation Metrics for Models Fine tuned on Imbalanced dataset
Model Precision Recall F1-Score
DeepSeek-7B 0.9241 0.9117 0.9110
Gemma-7B 0.9656 0.9451 0.9523
LLaMA-3.1-8B 0.9742 0.9730 0.9730
Mistral-7B 0.9715 0.9724 0.9720
CodeLLaMA-7B 0.9571 0.9533 0.9532
III-E Analysis of Impact of Context Window
To evaluate the effect of varying the context window, we experimented with two different window sizes: 200 and 300 characters surrounding the candidate secret string. These windows help the model understand the surrounding semantics, which is crucial for accurate classification.
Table XIV shows the F1-scores obtained by each fine-tuned model when tested under these two context window sizes.
TABLE XIV: Impact of Context Window Size on F1-Score (Fine-Tuned Models)
Model F1-Score (200 chars) F1-Score (300 chars)
DeepSeek-7B 0.9472 0.9690
Gemma-7B 0.9753 0.9764
LLaMA-3.1-8B 0.9852 0.9882
Mistral-7B 0.9819 0.9861
CodeLLaMA-7B 0.9707 0.9721
The results indicate that increasing the context window from 200 to 300 characters leads to a modest improvement in F1-score across all models. This improvement is attributed to the additional contextual cues available to the model for making more informed classifications—especially useful for complex or multiline secrets.
However, the gains are small, indicating that LLMs, even with smaller context windows, are already powerful enough to perform well due to their pretraining on massive and diverse code/text corpora. Importantly, smaller context windows reduce training and inference times, resulting in lower memory usage and faster model response—an important consideration for deployment in time-sensitive or resource-constrained environments.
Therefore, the choice of context window is a tradeoff between accuracy and resource, with 200 characters already offering strong performance in most use cases.
III-F Inference Time Analysis
We evaluated inference speed using 3,000 sequential predictions on a local GPU setup. The average inference time per sample was 0.1924 seconds on an RTX 4090, demonstrating that our models support real-time or near-real-time deployment in environments like IDEs or CI/CD pipelines.
III-G Resource Efficiency and Cost
All experiments ran on a workstation with an Intel i5-13400F CPU, 128 GB RAM, and an RTX 4090 GPU. Thanks to QLoRA and PEFT, we fine-tuned 7B–8B models efficiently within local hardware limits. This enables high-performing secret detection without costly cloud infrastructure, supporting privacy-focused and budget-conscious deployments.
IV Related Work
Large-Scale Secret Leak Analysis
Meli et al. [15] conducted a large-scale study on secret leaks within GitHub by analyzing billions of files through real-time scans and snapshots. They found that over 100,000 repositories expose thousands of secrets daily. Their regex, entropy-based filtering, and pattern detection approach achieved 99.29% accuracy.
Program Analysis for Secret Detection
Building upon earlier approaches, Sinha et al. [16] enhanced secret detection by incorporating program slicing, keyword-based, and pattern-based searches. Their method achieved high precision and recall for detecting API keys while also addressing false positives and the limitations of proximity-based detection.
Machine Learning-Based Filtering
To further reduce false positives, Saha et al. [17] introduced machine learning-based methods. They created a dataset of secrets and non-secrets from GitHub using regex and trained models on binary features like character patterns and entropy. Their Soft Voting Classifier achieved an F1-score of 86.7%.
Deep Learning Approaches
Feng et al. [19] explored deep learning techniques for detecting secrets, particularly textual passwords. By leveraging deep neural networks, they captured semantic patterns and significantly reduced false positives, demonstrating deep learning’s advantages over traditional rule-based methods.
Secret Detection in Issue Reports
Wahab et al. [21] focused on secret detection within issue reports and documentation using transformer-based language models such as BERT and RoBERTa. Their results highlighted the benefits of contextual embeddings in improving detection accuracy for non-code textual software artifacts.
Secret Mitigation Techniques
In terms of mitigation, Sinha et al. [40] suggested strategies like excluding secrets from version control, encrypting them using tools such as git-crypt, and integrating automated scanning mechanisms. They also emphasized practices like commit signing and two-factor authentication (2FA) to strengthen security.
Best Practices for Secret Management
Best practices for secret management were compiled by Basak et al. [23], who proposed 24 actionable guidelines. These included avoiding hard-coded secrets, using environment variables, employing short-lived credentials, utilizing secret managers like Vault, and sanitizing version control system (VCS) history.
Comparative Analysis of Detection Tools
Finally, Basak et al. [14] conducted a comparative analysis of both open-source and proprietary secret detection tools. They evaluated tools based on categorization, scoring methods, and detection capabilities, providing insights to help practitioners choose solutions based on precision, recall, and usability.
While prior work has predominantly focused on regex-based or machine learning-based techniques for detecting secrets in textual artifacts, our research shifts the focus to using large language models (LLMs) for identifying secrets directly within source code. By fine-tuning these models, we significantly reduce false positives and enhance detection accuracy, effectively bridging the gap between traditional approaches and modern deep learning within real-world development workflows.
V Threats to Validity
Internal validity concerns potential issues within the study that could affect the credibility of the results. An internal threat stems from hyperparameter tuning. While we adopted LoRA-based fine-tuning with reasonable settings, we did not perform exhaustive hyperparameter search due to resource constraints. Future experiments could explore further optimization to improve generalization. Additionally, since SecretBench [24] is derived from public GitHub repositories, there exists a potential risk of data contamination i.e., LLMs used in our study may have been exposed to similar or identical code during pretraining. However, we posit that even if some of this content was seen during training, the LLMs were unlikely to have been explicitly optimized to detect secrets in those contexts. As such, the learned knowledge is probably not aligned with the secret detection task, and our fine-tuning and prompt-based strategies remain essential for effective performance.
External validity refers to the generalizability of our results beyond the dataset and models used. Our evaluation focuses on a curated benchmark of GitHub repositories and fine-tuned models. These models may not generalize well to proprietary repositories or other languages and frameworks not represented in our dataset. To reduce this threat, we included diverse programming languages, secret types, and file types, aiming to simulate real-world scenarios. Nonetheless, further testing on industrial codebases would be necessary to validate broader applicability.
Construct validity addresses whether the study effectively captures what it aims to measure—in this case, true secret detection. One threat is the subjectivity in prompt design and few-shot examples used in LLM prompting. The selection of examples might influence model behavior, particularly for ambiguous candidates. We mitigated this by using representative, diverse examples and evaluating multiple prompting techniques (zero-, one-, and few-shot). Another construct-related concern is the exploitability of detected secrets. Although labeled as true secrets, not all exposed tokens may be exploitable or valid. Due to ethical and legal reasons, we did not test real-world exploitability. This limitation may affect interpretations of the model’s effectiveness in practice.
Conclusion validity reflects the accuracy of the interpretations made from our experimental results. To ensure robust conclusions, we used standard evaluation metrics—precision, recall, and F1-score—along with confusion matrix analysis. Still, model performance can vary across deployment scenarios, especially in noisy or highly imbalanced environments. We attempted to reduce individual evaluation bias by relying on automated classification and consistent test datasets. However, model-specific peculiarities or overfitting to the fine-tuning dataset could skew the results. Future replication studies can help verify our findings across alternate settings.
VI Conclusion
In this study, we investigated the use of large language models (LLMs) for detecting secrets in source code, addressing the limitations of traditional regex and entropy-based tools, which often suffer from high false positive rates. We introduced a hybrid approach that first uses regex-based candidate extraction and then applies LLM-based classification to filter out false positives. Our results show that fine-tuned models such as LLaMA achieve near-perfect performance in both binary and multiclass classification tasks, while pre-trained models like GPT-4o also perform well in few-shot settings, demonstrating the flexibility of LLMs in security-related applications.
To support our experiments, we curated and balanced a subset from the SecretBench dataset, extracted contextual code snippets, and designed structured prompts for training and inference. Our extensive evaluation across prompting strategies and model types highlights the strong performance of fine-tuned open-source models, which offer practical benefits in environments with privacy constraints or limited access to commercial APIs.
Looking forward, our approach can be extended by integrating it into CI/CD pipelines and IDEs for real-time secret leak prevention. Future work may also explore ensemble methods or uncertainty-based rejection mechanisms to further minimize or eliminate false negatives. Additionally, expanding detection to other artifacts such as issue reports and improving generalization through domain adaptation for proprietary and multilingual codebases remains an important direction. Furthermore, the development of more effective and adaptive prompting techniques for security tasks may further boost the performance of pre-trained models.